In the laboratory a coffee cup calorimeter, or constant pressure calorimeter, is frequently used to determine the specific heat of a solid, or to measure the energy of a solution phase reaction. A student heats 65.42 grams of tungsten to 97.61 °C and then drops it into a cup containing 81.69 grams of water at 22.57 °C. She measures the final temperature to be 24.38 °C The heat capacity of the calorimeter (sometimes referred to as the calorimeter constant) was determined in a separate experiment to be 1.89 1°C. Assuming that no heat is lost to the surroundings calculate the specific heat of tungsten. Specific Heat (W)- J gC.
Answer #1

Heat lost by tungsten = heat gained by water + heat gained by calorimeter

m x Cp x (T2-T1) = m' x Cp' x (T1-T2') + C x (T1 - T2')

65.42 g x Cp x (97.61 - 24.38)c = 81.69g x 4.184 J/gC x (24.38 - 22.57)c + 1.89 J/C x (24.38 - 22.57)c

4790.7066 x Cp = 618.642 + 3.421

Cp = 0.1298 J/gC